OpenStudy (anonymous):

A bullet is fired at 32 degrees with respect to the horizontal. If the speed at which the projectile leaves the gun is 300 m/s find the magnitude of the horizontal and the vertical components of the bullet's initial velocity .

OpenStudy (anonymous):

The initial velocity is 300 m/s on the hypotenuse of a triangle with an angle of 32 degrees above the horizontal. Draw a triangle and use this information along with the sin and cos functions to find the initial vertical and horizontal velocities. Do you know how to use sin and cos?

OpenStudy (anonymous):

Sin is opposite over hypotenuse Cos=adj/ hypotenuse . But how am I supposed to use them if I don't have another value ?

OpenStudy (anonymous):

Yes that is correct. The answer comes from this equation: sin(angle)=opposite/hypotenuse. So you solve for the opposite and get hypotenuse*sin(angle)=opposite and you can get the opposite side's value

OpenStudy (anonymous):

So to put it in your calculator put 300*sin(32) and that will give you your answer and similarly with cos to find the adjacent

OpenStudy (anonymous):

Ohkayyyy, I think I got it , So I have to set up like a cross multiplication type thing right? Sin 32/ 1=a/300 So, a=(sin32) (300)?

OpenStudy (anonymous):

No it isn't really cross multiplication. Look at this link and look at the first triangle on the right. http://en.wikipedia.org/wiki/Trigonometric_functions This is how you would set up your triangle in this problem with h=300, and the angle A=32. The functions you typed in for sin and cos are correct but remember when solving this you are taking the sin and cos of an angle to get opp/hyp or adj/hyp.

OpenStudy (anonymous):

Ill set up the eq's for you to see if its more clear. So you know length h and angle A. Now to find length b you would use the equation cos(A)=b/h and multiply both sides by h to get h*cos(A)=b. Now plug in your numbers and get 300*cos(32) and you get length b, or your horizontal initial velocity

OpenStudy (anonymous):

What is b in the equation? the Adjacent leg or the opposite leg?

OpenStudy (anonymous):

The adjacent leg

OpenStudy (anonymous):

And for the future, jsut remember: If you have a 2-dimensional vector 'v' with length 'a' and angle 'φ' between the vector and the X-axis, the vector always is:\[v = \left(\begin{matrix}a \cos(\phi) \\ a \sin(\phi)\end{matrix}\right)\]You'll need this all_the_time ;)

OpenStudy (anonymous):

b=cos(32)/300? And @MuH4hA Alright, thanks (:

OpenStudy (anonymous):

b=cos(32)*300. Remember you have to multiply the h in b/h=cos(32) to make it go to the other side

OpenStudy (anonymous):

FML. -_- I'm stupid. Okay, so 254.4?

OpenStudy (anonymous):

Yeah thats right. Now find the vertical

OpenStudy (anonymous):

Would that be the Pythagorean theorem?

OpenStudy (anonymous):

yes either that or use h*sin(angle)=a

OpenStudy (anonymous):

once you get the hang of it its easiest to just use the sin and cos functions because they are exactly the same except the sin and cos (if that makes sense)

OpenStudy (anonymous):

I'm getting huge numbers . a^2+64719.36=90000?

OpenStudy (anonymous):

Use the sin function

OpenStudy (anonymous):

158.99?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ahhhhhhh!!! :D thank you SO MUCH.

OpenStudy (anonymous):

No problem. Good luck